OpenStudy (anonymous):

Rewrite the expression so that there is no denominator: (tan2x)/(cscx+1)

OpenStudy (amoodarya):

do you mean tan 2x or tan ^2 x ?

OpenStudy (anonymous):

tan 2x :)

OpenStudy (amoodarya):

\[\frac{ \tan 2x }{ \csc x+1 }=\frac{ \tan 2x }{ \frac{ 1 }{\sin x }+1 }=\]

OpenStudy (amoodarya):

\[\frac{ \tan 2x }{ \frac{ 1 }{\sin x }+1 }=\frac{ \tan 2x }{ \frac{ 1+sinx }{\sin x }}=\]

OpenStudy (amoodarya):

did you try it ?

OpenStudy (anonymous):

yea :)

OpenStudy (amoodarya):

and what was your problem ?

OpenStudy (anonymous):

I'm not sure how to go on from there to get rid of the denominator...

OpenStudy (amoodarya):

show me your work

OpenStudy (anonymous):

I stopped there, and I don't know how to get rid of the denominator...

OpenStudy (amoodarya):

\[\frac{ \tan 2x }{ \csc x+1 }\times \frac{ \csc x-1 }{ \csc x-1 }= \frac{ \tan 2x (\csc x-1) }{ \csc^2 x-1 }=\]

OpenStudy (amoodarya):

\[\csc ^2 x -1=\cot ^2 x\]

OpenStudy (amoodarya):

do you fin out what would you do ?

OpenStudy (amoodarya):

\[ \frac{ \tan 2x (\csc x-1) }{ \cot^2 x }=\tan 2x (\csc x-1) \tan^2 x\]

OpenStudy (anonymous):

Thank you so much! :)